OpenStudy (egracer):

CAN SOMEONE TELL ME IF THESE ARE CORRECT!! ANWSER ARE ON THE BOTTOM Part A: Factor 3x2y2 − 2xy2 − 8y2. Show your work. Part B: Factor x2 + 10x + 25. Show your work. Part C: Factor x2 − 36. Show your work. -------------------- Part A-y^2(3x^2y^2-2x-8) Part B-(x+5)^2 Part C-(x+6)(x-6)

Nnesha (nnesha):

if you have 3 different typez of chocolates in the boxand u will take out all of 3 how many will left in the box ?

OpenStudy (egracer):

0

Nnesha (nnesha):

yes right so factor menas take out common stuff right so first question when you take out y^2 from first term how many y will left in the parentheses \[\huge\rm 3x^2y^2 − 2xy^2 − 8y^2\]

Nnesha (nnesha):

or in other words divide all 3 terms by common factor

Nnesha (nnesha):

\[ \huge\rm \frac{ 3x^2y^2 }{ y^2 }-\frac{ 2xy^2 }{ y^2}-\frac{ 8y^2 }{ y^2 }\]

OpenStudy (egracer):

So the y^2 would cancel out?

Nnesha (nnesha):

YEP!

Nnesha (nnesha):

so y^2(??????????????)

OpenStudy (egracer):

so it would be 3x^2-2x-8?

OpenStudy (jdoe0001):

btw Part B-(x+5)^2 \(\checkmark\) Part C-(x+6)(x-6) \(\checkmark\)

OpenStudy (egracer):

thank you

Nnesha (nnesha):

yep my pleasure

OpenStudy (egracer):

so the final answer would be 3x^2-2x-8

Nnesha (nnesha):

who kidnapped y^2 ??

OpenStudy (egracer):

y^2(3x^2-2x-8)

Nnesha (nnesha):

lO_Oks gO_Od

OpenStudy (egracer):

THANKKK YOUUU SOOO MUCHHHHH

OpenStudy (jdoe0001):

hmmm

Nnesha (nnesha):

now factor 3x^2 -2x-8

OpenStudy (egracer):

hmmmm combine like terms first right?

Nnesha (nnesha):

are there any *like* terms ?

OpenStudy (egracer):

nope there are none

Nnesha (nnesha):

so i would apply *Headphone* method find two number if you multiply them you should get product of AC and if you add or subtract them you should get middle term a=leading coefficient b=middle term c=constant term

Nnesha (nnesha):

\[\huge\rm \color{reD}{A}x^2-\color{blue}{B}x-\color{green}{C }\] \[\huge\rm \color{reD}{3}x^2-\color{blue}{2}x-\color{green}{8 }\]

OpenStudy (egracer):

so suntract them?

Nnesha (nnesha):

no what is a and c ?

OpenStudy (egracer):

3-8?

Nnesha (nnesha):

yep a =3 and c=-8 so multiply them ac =3 times -8 = -24 find two number if you multiply them you should get -24 and if you add or subtract them you should get -2

OpenStudy (egracer):

so the answer is y^2(3x+4)(x-2)

Nnesha (nnesha):

yep that's right ..........

OpenStudy (egracer):

thankk youu

Nnesha (nnesha):

my pleasure